# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v2

@@ 1.2 Mach’s Principle
Alexander Unzicker explored galaxy dynamics in relation to Mach’s Principle and the rotating bucket of water experiment, and proposed a basis for a MOND to solve flat rotation curves without considering Milgrom’s MOND [34]. For Unzicker, the rotation of a galaxy is the reason for gravitation, and a modification to the gravitational constant in function of size, mass, and angular momentum of the galaxy could explain flat rotation curves. He first proposes another interpretation of Sciama’s formulation substituting $c^{2}$ for $v^{2}$ so that $v^{2}=G\sum{m_{i}/r_{i}}$ and considering $v$ the asymptotic rotation velocity of the galaxy, but lacks an interpolating function through which it can be effectively applied to the solar system. He proposes a thought experiment: According to Mach, the system of two distant masses rotating around their center of mass in equilibrium of gravitational and centrifugal forces in an empty universe, since rotation between them is undefined due to absence of absolute space, must be equivalent to the system of those two distant masses not rotating and without gravity (only radial or relative velocities are measurable, and tangential velocities are impossible to measure instantaneously). The maximum possible angular velocity is limited by $w_{max}=c/r$ so that the maximum tangential speed is $v=c$ for one of the masses rotating around the other one, since the speed of light cannot be surpassed in any case. This luminal rotation in absence of background is a feature of introducing Sciama’s relationship in Newton’s law of gravity.
@@
Jaume Gine explored a Machian interpretation of the modified second law of motion for the simple interpolating function of Milgrom’s MOND based on the accelerated expansion of the universe [30], in which $a_{0}=GM_{u}/{R_{u}}^{2}=c^{2}/R_{u}$ through Sciama’s relationship is the acceleration that an experimental body feels induced by the rest of the matter of the universe in its inertial frame of reference, and the value of acceleration for which inertial and gravitational masses of a body can differ. Thus, the equivalence principle between inertial and gravitational masses is broken for accelerations smaller than $a_{0}$. In a following paper [31], Gine attempted to derive a phenomenological version of Milgrom’s MOND modification to Newton’s second law. He considers $a_{0}$ to be the acceleration at which the edge of the universe at distance $R_{u}$ goes away from a considered central inertial point due to the expansion of the universe. By relativizing acceleration considering the distant universe at rest, a new interpolating function similar to Milgrom’s MOND simple interpolating function is derived.
@@
Even though interesting ideas have been put forward relating MOND to Mach’s principle, an exact equivalence will be presented onwards, which should arise as an approximation in any nonlinear theory of modified inertia or gravity which attempts to explain dark matter effects in galaxies. As Milgrom stated in one of his conferences, “the only system that is strongly general relativistic and in the MOND regime is the Universe at large”.
@@ 2 Milgrom’s MOND from Mach’s Principle
Starting from the well-known form of Milgrom’s MOND standard interpolating function correction to Newtonian gravity $g_{N}=GM/r^{2}$ or to Newton’s second law $F=m_{i}g$, the “true” gravitational acceleration $g$ is
$g=\frac{G}{\mu\left(\frac{g}{a_{0}}\right)}\frac{M}{r^{2}}\ \ or\ \ F=\mu\left(\frac{g}{a_{0}}\right)m_{i}g$ (1)
$\mu\left(\frac{g}{a_{0}}\right)=\frac{1}{\sqrt{1+\left(\frac{a_{0}}{g}\right)^{2}}}\ \ or\ \ \mu\left(\frac{g}{a_{0}}\right)=\frac{1}{1+\frac{a_{0}}{g}}$ (2)
with $a_{0}$ being Milgrom’s MOND acceleration or gravitational field intensity constant (also known as Milgrom’s MOND scale constant), $M$ the active gravitational mass, $m_{i}$ the inertial mass, and $r$ the distance between the centers of mass of the active and passive gravitational masses, so that for the motion of a test particle in a gravitational field in the low acceleration or deep MOND regime, $\mu\left(g/a_{0}\right)\rightarrow g/a_{0}$ and $g=\sqrt{a_{0}\ g_{N}}$, so that with $g=v^{2}/r$, $v=\sqrt[4]{GMa_{0}}$ for achieving flat rotation curves in agreement with the Tully-Fisher law in a remarkable simple way with a single introduced free parameter.
Milgrom’s MOND acceleration or gravitational field intensity constant $a_{0}$ is usually interpreted as a fundamental constant in standard MOND theory, but possibly related to the cosmological parameter of the cosmological constant $a_{0}\sim c^{2}\sqrt{\Lambda}$ with $\Lambda=8\pi G\rho_{vac}/c^{2}$ and $\rho_{vac}$ the energy density of the cosmological constant or dark energy according to GR, which coincides in orders of magnitude (a better match can be done with $a_{0}=c^{2}\sqrt{\Lambda/3}$, although integers and pi are onwards omitted for simplicity). In order to express $a_{0}$ in terms of the mass and radius of the universe, $\rho_{vac}$ is substituted for the matter density of the universe $\rho_{vac}\sim\rho_{u}\sim M_{u}/{R_{u}}^{3}$ with $M_{u}$ the causally connected mass to the local system of study through the radius $R_{u}$ of the observable universe, a relationship which also coincides in orders of magnitude. Newton’s gravitational constant $G$ is substituted for Sciama’s interpretation of Mach’s principle $G\sim c^{2}R_{u}/M_{u}$, so that $a_{0}\sim c^{2}/R_{u}$, and by reverse engineering, $v\sim c\sqrt[4]{M/M_{u}}$ in the equivalent deep MOND regime and $g\sim c^{2}\sqrt{M/M_{u}}/r$. Inserting both $g$ and $a_{0}$ into the original interpolating function, it is now scale-invariant in mass and distance, and the speed of light cancels out, resulting in a Machian MOND formulation
$g=\frac{c^{2}}{\frac{M_{u}}{R_{u}}\ \mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)}\frac{M}{r^{2}}\ \ or\ \ \frac{M}{r^{2}}=\frac{M_{u}/R_{u}}{c^{2}}\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)g$ (3)
@@
Thus, a transformation (4) can be identified between the non-inertial reference frame due to gravity of a local system with a mass parameter $M$ and distance $r$, and a global reference frame of observable mass $M_{u}$ and size $R_{u}$ of the universe. The variable changed by this transformation is either the gravitational force (or equivalently, a variable gravitational constant), or the inertial law. Note that in (1), the interpolating function depends on true acceleration, while in Machian MOND depends on Newtonian acceleration.
@@
It is trivial that, since only equivalent substitutions for $G$ and $a_{0}$ according to Sciama’s interpretation of Mach’s principle which yield the same values are done (omitting integers and mathematical constants), and $M$ and $r$ take the same meaning as in Milgrom’s MOND, the resulting correction (3) and (4) for rotation curves of galaxies is equivalent to standard Milgrom’s MOND. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}<<1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow 1$ and Newtonian gravity is restored. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow\sqrt{\frac{M/r^{2}}{M_{u}/{R_{u}}^{2}}}$, $v=c\sqrt[4]{M/M_{u}}$, and the speed of rotating stars around galaxies far away from their galactic centers is determined by the relative active gravitational masses between the local system and the global system. The resulting velocities in this case do not strongly depend on the values of $M_{u}$ due to the fourth root, so they are not too sensitive to the uncertainties of the measured mass value.
@@
It is also trivial that, integrating an average mass density of the homogeneous and isotropic universe with uniform matter density at large enough scales over the causally connected spherical volume, the integral results in $NM_{u}/{R_{u}}^{2}$ (with $N$ being a rational factor that can be omitted for simplicity since it does not affect the orders of magnitude of the relationship between $M_{u}$ and $R_{u}$). Thus, under Machian MOND, the numerator $M_{u}/{R_{u}}^{2}$ in (4) is the total field intensity (the gravitational field intensity without the gravitational constant $G$) of all masses of the universe at the point in which the correction is applied (omitting integers for simplicity), taking the same meaning as $M/{r}^{2}$ in the denominator, which is the field intensity of the masses of the local system at the point in which the correction is applied.
@@
By decreasing the mass of the universe within the observable radius $M_{u}$, the correction term (4) decreases the gravitational force. But, considering also Sciama’s relationship, the gravitational force increases (having the same observational consequence as considering Sciama’s relationship alone, the decrease of inertia), and local velocities increase. Decreasing the mass of the universe until the absence of any background ($M_{u}\rightarrow M,R_{u}\rightarrow r$), the transformation $\mu$ (4) reduces to some simple number $K$ which can be set to unity with the appropriate choice of integers in the formulation of the Machian transformation and Sciama’s Mach’s principle interpretation (a simple way to obtain $K=1$ is considering Schrödinger’s original definition of $G$ and multiplying the term in parenthesis of the transformation by $3$), and velocities approach $v\rightarrow Kc$ at the edge of the galaxy. This is already a feature of Sciama’s relationship, and it is reasonable according to Mach, since without any background, rotation is undefined and tangential speed can take up any value up to the limit of the speed of light. The proposed transformation respects this result and also respects the scale invariance in mass and distance that Sciama’s relationship introduces (doubling the mass or size of both the universe and the local system results in the same estimated velocities) to the scale-dependent Newtonian gravity (Milgrom’s MOND is only scale-invariant when $a_{0}\rightarrow\infty$ fixing $Ga_{0}$).
@@ 3 Discussion
It is shown that Milgrom’s MOND correction to Newtonian gravity (or inertia and Newton’s second law) can be reformulated based on Mach’s principle without dimensional constants or free parameters except for the speed of light, ensuring scale invariance.
@@
The relationship pointed out by Milgrom between his MOND’s acceleration constant $a_{0}$ and the cosmological constant or the Hubble parameters is interpreted merely as accidental due to the cosmological coincidence by which, at present epoch, the energy densities associated with dark energy (the cosmological constant) and visible matter are of the same orders of magnitude.
@@
The Machian MOND’s interpolating function is a transformation $\mu(M_{u},R_{u},M,r)$ (4) between different scales and non-inertial reference frames which, by using the standard MOND interpolating function, resembles a Lorentz-like transformation factor with a plus sign instead of a minus sign: the relationship between a local frame in spacetime, and that of a global frame which is at rest (the background of the rest of the universe) with respect to the first one, with both being gravitationally accelerated relative to each other. It expresses how gravity changes locally due to the background universe through relative masses and distances in both systems, in accordance with relational mechanics. The speed of light in Sciama’s relationship takes the meaning of the speed of gravity and causality justifying the choice of $R_{u}$ as the radius of the observable universe containing mass $M_{u}$, which is the part of the universe causally affecting local dynamics, for Mach’s interpretation not to violate locality.
@@
Machian MOND satisfies several definitions of Mach’s principle: Newton’s gravitational constant $G$ is a dynamical field, local inertial frames are affected by the cosmic distribution of matter, and rotation is undefined up to the speed of light in the absence of the rest of the universe. If the correction applies to inertia instead of just to gravity, then inertial mass would also be affected by the global distribution of matter and an isolated body in otherwise empty space would have no inertia. In agreement with Mach’s principle, if inertial mass is defined by the potential of the distribution of mass (following Sciama’s argument), it is reasonable to think that it could also be affected by the field intensity of the distribution of mass. Thus, Machian MOND introduces an even stronger (and nonlinear) dependence of local dynamics on the distribution of matter than theories that just consider Sciama’s relationship, and can be considered more Machian.
Milgrom himself favors modifying inertia (which is the true aim of Mach’s principle), under the justification that Newtonian inertia was already modified by special relativity. The formal limit where $a_{0}\rightarrow 0$ in which Milgrom’s MOND approaches the standard Newtonian action does not hold, since the mass and radius of the universe (for which $a_{0}$ is replaced) always exist. Machian MOND, or a similar and equivalent approximate formulation, is considered to be a necessary nonlinear feature of a phenomenological theory of modified inertia or modified gravity which incorporates Mach’s principle, since Milgrom’s MOND is clearly nonlinear due to its external field effect. This might arise by imposing Mach’s principle (for instance, through the relativity of accelerations and inertia, and inertia being of gravitational origin) to a non-Machian nonlinear theory of gravity, such as GR. So far, only linear and non-relativistic full Machian theories of modified inertia with Sciama’s relationship have been constructed.
Milgrom’s MOND originally violates conservation of linear momentum for N-body systems. This can be fixed through a non-relativistic Lagrangian formulation of MOND as a modification to Newtonian gravity, which satisfies all standard conservation laws [35]. Applying the Gauss law for cases of high symmetry, (1) is recovered. Machian MOND (3) can be derived from an action in the same way. As modified gravity, it does not violate Einstein’s equivalence principle (although the strong equivalence principle is violated) because the modification to the gravitational acceleration does not depend directly on the passive gravitational or inertial mass.
@@
It is well known that, according to Bekenstein, a higher value for Milgrom’s MOND $a_{0}$ could resolve the tension between standard MOND and galaxy cluster’s core dynamics (and that $a_{0}$ seems to grow with scale), where Newtonian accelerations take around that same value (provided that $a_{0}$ stays the same in galaxies for agreement with rotation curves). Machian MOND, although being equivalent to Milgrom’s MOND, is effectively a theory of a varying $a_{0}$ with time scale and could in principle resolve this issue of Milgrom’s MOND. It could also in principle solve the cosmological issues of Milgrom’s MOND [36]. As modified gravity, the equivalent variation of the Newtonian $G$ is both temporal due to the expanding universe and dependence on the mass and radius of the observable universe, and spatial since local potentials should also be included in Sciama’s relationship and the correction through the interpolating function depends on the local field intensity.
@@
It would be interesting to compare Machian MOND with observational constraints of a varying gravitational constant at the early universe considering a cosmological model without dark matter (assuming that Machian MOND solves partly the need for dark matter), due to its dependence on the mass and radius of the observable universe, which have both varied over time. A similar study has been done for quantized inertia by comparing the prediction of a specific increase in the galaxy rotation anomaly at higher redshifts [37]. The formulation of quantized inertia is similar to the one proposed in this article (without Sciama’s relationship), but its interpretation is not based on Mach’s principle. In fact, no Planck’s constant is used in the quantized inertia formulation, which suggests that its origin is not related to quantum phenomena.

### v3

@@ 1.2 Mach’s Principle and modified inertia
Alexander Unzicker explored galaxy dynamics in relation to Mach’s Principle and the rotating bucket of water experiment, and proposed a basis for a MOND to solve flat rotation curves without considering Milgrom’s MOND [26]. He proposes the following thought experiment. According to Mach, the system of two distant masses rotating around their center of mass in equilibrium of gravitational and centrifugal forces in an empty universe, since rotation between them is undefined due to absence of absolute space, must be equivalent to the system of those two distant masses not rotating and without gravity (only radial or relative velocities are measurable, and tangential velocities are impossible to measure instantaneously, as they are indistinguishable from a rotation of the coordinate system of the observer). The maximum possible angular velocity is limited by $w_{max}=c/r$ so that the maximum tangential speed is $v=c$ for one of the masses rotating around the other one, since the speed of light cannot be surpassed in any case. This luminal rotation in absence of a background universe from Unzicker’s thought experiment is a feature of introducing Sciama’s relationship in Newton’s law of gravity, in which the Newtonian $v=\sqrt{GM/r}$ changes to $v=c\sqrt{(M/r)/(M_{u}/R_{u})}$ after introducing $G=c^{2}/(M_{u}/R_{u})$ and without a background universe $M_{u}\rightarrow M$ and $R_{u}\rightarrow r$, velocities are $v=c$.
@@
Jaume Gine explored a Machian interpretation of the modified second law of motion for the simple interpolating function of Milgrom’s MOND based on the accelerated expansion of the universe [30], in which $a_{0}=GM_{u}/{R_{u}}^{2}=c^{2}/R_{u}$ through Sciama’s relationship is the acceleration that an experimental body feels induced by the rest of the matter of the universe in its inertial frame of reference, and the value of acceleration for which inertial and gravitational masses of a body can differ. Thus, the equivalence principle between inertial and gravitational masses is broken for accelerations smaller than $a_{0}$. In a following paper [31], Gine attempted to derive a phenomenological version of Milgrom’s MOND modification to Newton’s second law. He considers $a_{0}$ to be the acceleration at which the edge of the universe at distance $R_{u}$ goes away from a considered central inertial point due to the expansion of the universe. By relativizing acceleration considering the distant universe at rest, he derives a new interpolating function similar to Milgrom’s MOND simple interpolating function.
@@
Even though interesting ideas have been put forward relating MOND to Mach’s principle, an exact equivalence and its implications will be presented onwards, which should arise as an approximation in any non-linear theory of modified inertia or gravity which attempts to explain dark matter effects in galaxies. The motivation behind this interpretation of Milgrom’s MOND is that it can be reformulated with the same parameters of the mass and radius of the observable and causally connected universe that the Machian models of modified inertia introduce.
@@ 2 Milgrom’s MOND from Mach’s Principle
Starting from the well-known form of Milgrom’s MOND standard or simple interpolating function correction to Newtonian gravity $g_{N}=GM/r^{2}$ or to inertia $F=m_{i}g$, the ‘true’ gravitational acceleration $g$ or ‘true’ second law of inertia is
$g=\frac{G}{\mu\left(\frac{g_{N}}{a_{0}}\right)}\frac{M}{r^{2}}\ \ or\ \ F=\mu\left(\frac{g}{a_{0}}\right)m_{i}g$ (1)
$\mu\left(\frac{a}{a_{0}}\right)=\frac{1}{\sqrt{1+\left(\frac{a_{0}}{a}\right)^{2}}}\ \ or\ \ \mu\left(\frac{a}{a_{0}}\right)=\frac{1}{1+\frac{a_{0}}{a}}$ (2)
with $a_{0}$ being Milgrom’s MOND acceleration scale constant, $g_{N}=GM/r^{2}$ the Newtonian gravitational acceleration, $M$ the active gravitational mass, $m_{i}$ the inertial mass, and $r$ the distance between the centers of mass of the active and passive gravitational masses, so that for the motion of a test particle in a gravitational field in the low acceleration or deep MOND regime, $\mu\left(a/a_{0}\right)\rightarrow a/a_{0}$ and $g=\sqrt{a_{0}\ a}$, so that with $g=v^{2}/r$, $v=\sqrt[4]{GMa_{0}}$ for achieving flat galaxy rotation curves in agreement with the Tully-Fisher law in a remarkable simple way with a single new free parameter.
Milgrom’s MOND acceleration scale constant $a_{0}$ is usually interpreted as a fundamental constant in MOND theory, but possibly related to the cosmological parameter of the cosmological constant $a_{0}\sim c^{2}\sqrt{\Lambda}$ or the energy density of the vacuum $\rho_{vac}$ with $\Lambda=8\pi G\rho_{vac}/c^{2}$ according to GR, which coincides in orders of magnitude (a better match can be done with $a_{0}=c^{2}\sqrt{\Lambda/3}$, although integers and Pi are onwards omitted for simplicity). This is due to Milgrom’s wish to define absolute acceleration with respect to the vacuum energy frame in MOND as modified inertia. But Mach’s principle implies that acceleration should be defined with respect to the rest of the masses of the universe. In order to express $a_{0}$ in terms of the mass and radius of the universe, $\rho_{vac}$ is substituted for the matter density of the universe $\rho_{vac}\sim\rho_{u}\sim M_{u}/{R_{u}}^{3}$ with $M_{u}$ the causally connected mass to the local system of study through the radius $R_{u}$ of the observable universe (or particle horizon radius), a relationship which also coincides in orders of magnitude. Newton’s gravitational constant $G$ is substituted for Sciama’s interpretation of Mach’s principle $G\sim c^{2}/(M_{u}/R_{u})$, so that $a_{0}\sim GM_{u}/{R_{u}}^{2}\sim c^{2}/R_{u}$, and by reverse engineering, $v\sim c\sqrt[4]{M/M_{u}}$ and $g\sim c^{2}\sqrt{M/M_{u}}/r$ in the equivalent deep MOND regime. The interpolating function is now scale-invariant in mass and length, resulting in a Machian MOND formulation
$g=\frac{c^{2}}{\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\ \mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)}\frac{M}{r^{2}}\ \ or\ \ \frac{M}{r^{2}}=\frac{\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)}{c^{2}}\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)g$ (3)
@@
in which for most cases where local potentials are small, $\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\sim\left(\frac{M_{u}}{R_{u}}\right)$. Inertial mass is then defined as the scalar $m_{i}=m_{g}G(M_{u}/R_{u})/c^{2}$ having an origin in the gravitation of all other masses in the universe, with $m_{g}$ the gravitational mass.
@@
It is trivial that, since only equivalent substitutions for $G$ and $a_{0}$ according to Sciama’s interpretation of Mach’s principle which yield around the same values are done (omitting integers and mathematical constants), and $M$ and $r$ take the same meaning as in Milgrom’s MOND, the resulting correction (3) and (4) for galaxy rotation curves are equivalent to Milgrom’s MOND. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}<<1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow 1$ and equivalent Newtonian laws are restored with $v=c\sqrt{({M/r})/({M_{u}/R_{u}})}$. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow\sqrt{\frac{M/r^{2}}{M_{u}/{R_{u}}^{2}}}$, $v=c\sqrt[4]{M/M_{u}}$, and the speed of rotating stars around galaxies far away from their galactic centers is determined by the relative active gravitational masses between the local and the global system.
@@
It is also trivial that integrating an average mass density of the homogeneous and isotropic universe at large enough scales over the causally connected spherical volume results in $NM_{u}/{R_{u}}^{2}$ (with $N$ being a dimensionless factor that can be omitted for simplicity, as it does not affect the orders of magnitude of the relationship between $M_{u}$ and $R_{u}$). A precise calculation would require considering Lienard-Wiechert potentials and the retarded Green functions to account for the propagation of gravity at the speed of light and current mass density instead of simple mass density. Under Machian MOND, the denominator $M_{u}/{R_{u}}^{2}$ in (4) is the total field intensity (the gravitational field intensity without the gravitational constant $G$) of all masses of the universe (omitting dimensionless factors for simplicity), taking the same physical meaning as $M/{r}^{2}$ in the numerator, which is the field intensity of the masses of the local system at the point in which the correction is applied.
Thus, a transformation (4) can be identified between the non-inertial reference frame of a local system with a mass parameter $M$ and distance $r$, and a global reference frame of observable mass $M_{u}$ and size $R_{u}$ of the universe. The variable changed by this transformation is either the gravitational force (equivalent to a variable gravitational constant) or the inertial law. Note that in (1), the interpolating function in modified inertia depends on true acceleration, while in Machian MOND it depends on the local field intensity, which suggests that inertia has an origin in gravitation, as the modified inertia theories of Sciama and Treder suggest through the same parameters $M_{u}$ and $R_{u}$ present in Machian MOND.
@@
By decreasing the mass of the universe within the observable radius, the transformation (4) decreases the gravitational force when $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$. But, considering also Sciama’s relationship for the gravitational constant, the gravitational force increases (having the same observational consequence as considering Sciama’s relationship alone, the decrease of inertia), and local velocities increase. The formal limit where $a_{0}\rightarrow 0$ in which Milgrom’s MOND approaches Newton’s laws does not hold, since the mass and size of the universe (for which $a_{0}$ is effectively replaced) always exist. The mass of the universe can never be zero in the system under study, it can only be small as the very same mass of the system itself. Decreasing the mass of the universe until the absence of any background ($M_{u}\rightarrow M,R_{u}\rightarrow r$), the transformation (4) reduces to some simple number $K$ which can be set to unity with the appropriate choice of integers in the formulation of the Machian transformation and Sciama’s Mach’s principle interpretation (a simple way to obtain $K=1$ is considering Schrödinger’s original definition of $G$ and multiplying the term in brackets of the transformation by $3$), and velocities approach $v\rightarrow Kc$. This is already a Machian feature of Sciama’s relationship by which without any background, rotation is undefined, and tangential speed can take any value up to the limit of the speed of light. The proposed transformation respects this result and also respects the scale invariance in mass, length, and time that Sciama’s relationship introduces in the Newtonian regime (classical Newtonian gravity is scale-dependent). Milgrom’s MOND is only length and time scale-invariant when $a_{0}\rightarrow\infty$ fixing $Ga_{0}$ to achieve constant velocities independent of radius at the flat part of the galaxy rotation curves. Machian MOND is thus scale-invariant in mass, length, and time $(m,r,t)\rightarrow(\lambda m,\lambda r,\lambda t)$ at all regimes. Doubling the mass or size of both the universe and the local system always results in the same estimated velocities.
@@ 3 Discussion
It is shown that Milgrom’s MOND correction to Newtonian gravity or inertia can be reformulated based on Mach’s principle without dimensional constants or free parameters except for the speed of light, ensuring mass, space and time scale invariance at all regimes.
@@
The relationship pointed out by Milgrom between his MOND’s acceleration constant $a_{0}$ and the cosmological constant or the Hubble parameter is interpreted merely as accidental due to the cosmological coincidence by which, at the present epoch, the energy densities associated with dark energy (the cosmological constant) and visible matter are of the same orders of magnitude.
@@
The Machian MOND’s transformation is a function $\mu(M_{u},R_{u},M,r)$ (4) between different scales and non-inertial reference frames: the relationship between a local frame in spacetime, and that of a global frame which is at rest (the background of the rest of the universe) with respect to the first one, with both being accelerated relative to each other. It expresses how gravity or inertia changes locally due to the background universe through relative masses and distances in both systems, in accordance with relational mechanics, and respects the weak equivalence principle since it does not depend on the inertial mass. The speed of light in Sciama’s relationship takes the meaning of the speed of gravity and causality justifying the choice of $R_{u}$ as the radius of the observable universe containing mass $M_{u}$, which is the part of the universe causally affecting local dynamics, for Mach’s interpretation not to violate locality.
Although the origin of the Machian MOND transformation is not explained or derived here, it seems that it follows from Mach’s principle’s argument about which frame is acceleration defined with respect to. When the local field intensity is above that of the rest of the universe (at the cores of galaxies or within the solar system), the ‘main frame’ acceleration is defined with respect to is the local one, and one can describe the dynamics of the system without referring to the field intensity of the universe at large through Newtonian laws. But, if the field intensity of the rest of the universe is greater than the one from the local system (in the deep MOND regime, such as the flat velocity part of the disk in galaxies), it seems that acceleration is defined with respect to the background frame of the universe, and the field intensity of the universe at large must be taken into account non-linearly together with that of the local frame, to explain the dynamics of the local system (in the case of Machian MOND, effectively through the Machian MOND transformation (4) ). In between both regimes, the identification of the ‘main frame’ is shared between the local and global frames, described approximately by the slope of the Machian MOND transformation.
@@
Machian MOND satisfies several definitions of Mach’s principle: Newton’s gravitational constant $G$ is a dynamical field, local inertial frames are affected by the cosmic distribution of matter, and rotation is undefined up to the speed of light in the absence of the rest of the universe. If the correction applies to inertia instead of just gravity, then inertial mass is also affected by the global distribution of matter, and an isolated body in otherwise empty space would have no inertia. In agreement with Mach’s principle, if inertial mass is defined by the potential of the mass distribution (following Sciama’s argument), it is reasonable to think that it could also be affected by the field intensity of the mass distribution. Thus, Machian MOND introduces an even stronger (and non-linear, due to the external field effect of MOND) dependence of local dynamics on the distribution of matter than linear theories that just consider Sciama’s relationship, and it can be considered more Machian.
Milgrom himself favors modifying inertia (which is the true aim of Mach’s principle), and not just gravity. An example of modified inertia is that of SR $F=m_{i}d(\gamma v)/dt$ with $\gamma$ being the Lorentz factor [34]. The original Milgrom’s MOND as modified inertia with the inertial term of the equation of motion $\mu\left(g/a_{0}\right)m_{i}g$ is translation, rotation, and Galilei invariant, but leads to non-conservation of momentum in a many-body systems because the equation of motion is not derivable from an action. As a modification of the kinetic term in the Lagrangian, it conserves momentum, energy, and angular momentum, but is not Galilei invariant. A non-relativistic, MOND, kinetic action for a particle with conservation laws of momentum, energy, and angular momentum, and the boost symmetry being Galilean, must be time non-local [10]. But this only occurs if it is a function of the trajectory $r(t)$ (since the function requires knowledge of the entire trajectory) and it is required to have a MOND limit when $a_{0}\rightarrow 0$, which is neither the case for Machian MOND. We have omitted the original negative sign of Sciama’s relationship $c^{2}\sim-GM_{u}/R_{u}$ for simplicity. But by considering it and changing the sign in Milgrom’s MOND standard interpolating function from positive to negative for consistency, the resulting transformation
$\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{\sqrt{1-\left(\frac{-M_{u}/{R_{u}}^{2}}{M/r^{2}}\right)}}$ (5)
is equivalent to Machian MOND (4) and resembles the structure of a Lorentz-like transformation $\gamma=1/\sqrt{1-\beta}$. Another possibility is that this negative sign comes from $M/r^{2}$ having an opposite direction to that of $M_{u}/R_{u}^{2}$. If a function depending on velocities, such as the Lorentz factor, arises from Lorentz invariance in SR, where accelerations are absolute for consistency of the theory, one would expect a function depending on accelerations or field intensities, such as Machian MOND, to arise when relativizing accelerations. Lorentz invariance implies that the speed of light is constant in all inertial frames, and it seems that Machian MOND implies that the total field intensity at any point in the universe is always $>M_{u}/R_{u}^{2}$. It may be that Machian MOND arises from a new boost symmetry generalizing Lorentz invariance, for instance, through the relativity of acceleration, which is broken by the presence of the universe at large. Without a background universe, this symmetry would reduce to the Galilean or Lorentz one. If inertia has an origin in gravitation, then the modification on the kinetic term can be based only on the field intensity, without modifying the Poisson equation (and leaving gravity intact), which is shown in Machian MOND as modified inertia. The resulting Machian MOND formulation can be thought of as an in-between interpretation of modified inertia and modified gravity.
Machian MOND, or a similar and equivalent approximate formulation, is considered to be a necessary non-linear feature that a phenomenological theory of modified inertia or modified gravity which incorporates Mach’s principle reduces to as an approximation, in order to agree with galaxy rotation dynamics. This might arise by imposing Mach’s principle (for instance, through the relativity of accelerations and inertia of local objects being determined by the sources of the gravitational field at the global scale) to a non-Machian non-linear Lagrangian or to a theory of gravity such as GR (in which solutions to local dynamics do not depend on the mass or size of the observable universe). In fact, analogous interpolating functions in SR and GR have, as independent variables, dimensionless ratios of the form $v/c$ and $MG/Rc^{2}$, similar to Machian MOND. The transformation function should arise naturally as an approximation, without the need of introducing it by hand (as in AQUAL or QUMOND), but the global $M_{u}$ and $R_{u}$ terms governing local dynamics should appear in the complete FundaMOND theory, which is not the case in GR. So far, only linear and non-Lorentz invariant Machian theories of modified inertia with Sciama’s relationship have been constructed, such as that of Hans-Jürgen Treder [21].
@@
It is well known that, according to Bekenstein, a higher value for Milgrom’s MOND $a_{0}$ could resolve the tension between Milgrom’s MOND and galaxy cluster’s core dynamics (and that $a_{0}$ seems to grow with scale), where Newtonian accelerations take around that same value (provided that $a_{0}$ stays the same in galaxies for agreement with rotation curves). Machian MOND, although being equivalent to Milgrom’s MOND, is effectively a model of a varying $a_{0}$ with time scale and could in principle help resolving this issue of Milgrom’s MOND. It could also in principle solve the cosmological issues of Milgrom’s MOND [35]. As modified gravity, the equivalent variation of the Newtonian $G$ is both temporal due to the expanding universe and dependence on the mass and radius of the observable universe, and spatial since local potentials and local field intensities should also be included in Sciama’s relationship and the Machian MOND transformation, respectively.
@@
It would be interesting to compare Machian MOND with observational constraints of a varying effective gravitational constant at the early universe considering a cosmological model without most or all physical dark matter (assuming that Machian MOND solves at least partly the need for dark matter), due to its dependence on the mass and radius of the observable universe, which have both varied over time. A similar study has been done for quantized inertia by showing the prediction of a specific increase in the galaxy rotation anomaly at higher redshifts [36], but quantized inertia only makes the $a_{0}$ of Milgrom’s MOND dependent on $R_{u}$ through $a_{o}\sim c^{2}/R_{u}$. The formulation of quantized inertia is similar to the one proposed (without Sciama’s relationship), but its interpretation is not based on Mach’s principle. In fact, no Planck constant appears in the formulation, which suggests that its origin is not related to quantum phenomena, while the need to account for the universe at large directly points towards Mach’s principle. More recently, it has been observed that there is no significant evolution in the baryonic Tully-Fisher law for redshifts up to $z\sim 2$ [37], implying an almost constant $a_{o}$ in MOND. In contrast, Machian MOND shows a difference dependence compared to quantized inertia: the Machian MOND function depends not only on $R_{u}$, but also on $M_{u}$. If the term $M_{u}/R_{u}^{2}$ is more constant than the simple dependence on $R_{u}$, this observation would not rule out Machian MOND (the observable radius was smaller in the past, but so was the mass contained in the volume defined by the observable radius). But as explained before, Machian MOND depends on many uncertainties, such as the Hubble tension, whether some physical dark matter is present in $M_{u}$, missing integers and Pi in the formulation, and relativistic effects of the universe as a whole.